# Compressing Transformer-Based Self-Supervised Models for Speech Processing

### v2

@@ Abstract
Despite the success of Transformers in self-supervised learning with applications to various downstream tasks, the computational cost of training and inference remains a major challenge for applying these models to a wide spectrum of devices. Several isolated attempts have been made to compress Transformers, but the settings and metrics are different across studies. Trade-off at various compression rates are also largely missing in prior work, making it difficult to compare compression techniques. In this work, we aim to provide context for the isolated results, studying several commonly used compression techniques, including weight pruning, head pruning, low-rank approximation, and knowledge distillation. We report trade-off at various compression rate, including wall-clock time, the number of parameters, and the number of multiply-accumulate operations. Our results show that compared to recent approaches, basic compression techniques are strong baselines. We further present several applications of our results, revealing properties of Transformers, such as the significance of diagonal attention heads. In addition, our results lead to a simple combination of compression techniques that improves trade-off over recent approaches. We hope the results would promote more diverse comparisons among model compression techniques and promote the use of model compression as a tool for analyzing models. Our code of compressing speech self-supervised model is available at https://github.com/nervjack2/Speech-SSL-Compression/.
@@
Compression, Self-supervised, Representation
@@ I Introduction
Self-supervised learning for speech processing has had a great success in scaling up with the increasing amount of unlabeled data, achieving strong performance in various tasks, such as automatic speech recognition (ASR), and speaker verification [1]. To make use of the large amount of unlabeled data, self-supervised models are typically large, for example, with 12 or more layers of Transformers [2, 3, 4, 5]. These large models are slow in runtime, requiring cutting-edge hardware, such as GPUs with sufficient memory. In this work, we focus on compressing Transformer-based self-supervised models, making them more efficient and suitable for a wide spectrum of devices and settings.
Common compression techniques, such as iterative pruning [6, 7, 8], low-rank approximation [9, 10, 11], and knowledge distillation [12], have little in common—a sign that the concept of model compression is broad, if not too broad. Model compression is typically defined as a function $C$ that takes in a model $m$ and returns a small model $C(m)$ (sometimes with additional training and data). The compression function $C$ can, in fact, return a small model, completely ignoring the input model $m$. For example, training a small model from random initialization and neural architecture search can both be treated as model compression, even though nothing is really compressed. Any approach that returns a small yet performant model contributes to the trade-off and should be considered in the opportunity cost of compression. In this work, we limit the study to iterative pruning, low-rank approximation, and knowledge distillation. Following this argument, we also provide simple baselines, where we simply take the early layers, i.e., stopping the forward process of the large model early. Stopping the forward process early not only reduces computation but also maintains satisfactory performance, because low layers are often already useful for several downstream tasks [13].
Comparing compression techniques involving self-supervised training is complicated. Some study compression for a particular downstream task [14, 15, 16, 17], while some aim to obtain a model that can extract a general representation across different downstream tasks [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In addition to the common task accuracy, some compare the number of parameters [19], while some compare wall-clock time [21, 25] or the number of multiply-accumulate operations (MACs) [17]. Many have also failed to include simple baselines, making it difficult to compare approaches.
In this work, we focus on compressing self-supervised models with the self-supervised loss that the models are trained with. As a baseline, we choose MelHuBERT [28], a variant of HuBERT that takes Mel spectrograms as input and has a simplified training recipe. We train our own MelHuBERT from random initialization to obtain a complete training loss curve. Having a complete loss curve is essential because subsequent compression all rely on monitoring and maintaining the loss values.
To facilitate the comparison of compression techniques on self-supervised models, we provide a comprehensive study of theoretical and practical speed-up of different compression techniques, measuring wall-clock time, numbers of parameters, and numbers of MACs per one second speech, charting the landscape of compressing Transformer-based self-supervised models. Our results enable a fair and comprehensive comparison among prior work, such as DistilHuBERT [19], FitHuBERT [21], and LightHuBERT [20]. We also show that our simple baselines are strong contenders for these more involved approaches.
Our results suggest that pruning entire layers is the best for improving the runtime on GPUs, while head pruning and low-rank approximation are useful for reducing the number of parameters but not as effective in improving the runtime on GPUs. The results motivate us to chain these techniques together, leveraging the strengths of each technique. These simple combinations provide better trade-off than prior work.
Finally, we conduct an in-depth analysis on the compression results. We attempt to answer to what degree we can maintain the self-supervised loss when compressing and to what degree maintaining the self-supervised loss leads to generalization of different downstream tasks. In addition, prior work compress various parts of a neural network without providing insights as to what are compressed. We will show what parts tend to be compressed more, e.g., what types of attention heads tend to be compressed. Compression, in general, is a valuable tool for understanding models, and, in our case, Transformers.
We summarize our contributions as follows.
  1. 1.
    Different compression techniques have different strengths among MACs, numbers of parameters and real-time factors. We hope that this work will encourage others to evaluate compression methods from a more diverse and comprehensive perspective.
  2. 2.
    Mapping comprehensive trade-off allows us to compare several recent approaches, namely, DistilHuBERT [19], FitHuBERT [21], and LightHuBERT [20]. Our simple baselines are surprisingly strong compared to prior work.
  3. 3.
    We show that it is possible to provide better trade-off by simply chaining several compression techniques together.
  4. 4.
    We analyze what parts of Transformers are likely to be pruned, revealing various properties of self-supervised Transformers.
II Model Compression
Since the training loss after compression typically becomes worse, it is common to have additional training on the compressed model to recover the training loss. To monitor and to assess how much training loss is recovered, we need to have access to the training curve prior to any compression.
We choose MelHuBERT [28] due to its strong performance and simple training recipe. The simple training recipe allows us to obtain the entire training curve from random initialization. We will briefly review MelHuBERT and its self-supervised loss function, as the loss function will be what we monitor and want to recover.
After reviewing MelHuBERT, we will detail the specifics of model compression techniques used in this work, including weight pruning, head pruning, low-rank approximation, knowledge distillation. Except knowledge distillation, the others can be seen as instances of iterative pruning. Iterative pruning performs the following two steps iteratively.
  1. 1.
    Prune a block of weights.
  2. 2.
    Train the pruned network until the loss (to be introduced next) converges.
What constitutes a block and how a block is chosen can be customized. If we choose individual weights as blocks, this algorithm becomes weight pruning [29]; If we instead choose attention heads as blocks, this algorithm becomes head pruning [8]; If we choose rows or columns as blocks, this algorithm becomes low-rank approximation [9]. As opposed to one-shot pruning, we perform pruning iteratively, because iterative pruning typically achieves better compression [30]. We avoid using more involved variants of pruning as our baselines, such as those in [17, 25], because pruning while training or additional terms in the optimization add confounding factors, complicating our analysis.
II-A MelHuBERT
MelHuBERT [28] is a variant of HuBERT, with simplifications on both the model architecture and the training objective. MelHuBERT takes Mel spectrograms as input (hence the name), as opposed to wave samples in HuBERT. MelHuBERT excludes the seven convolution layers in HuBERT for producing 20-ms frames, saving about 33.3% of MACs and about 31.2% of pre-training time on a single 24GB RTX 3090 GPU. In addition to the model architecture, the training objective of MelHuBERT has been simplified as well. Instead of using the wav2vec 2.0 style loss function in HuBERT, MelHuBERT minimizes the cross entropy loss, which has been shown to accelerate training [31].
Since the loss will be used for subsequent compression, we formally define it as follows. Suppose we have a sequence of 10-ms log Mel frames $x_{1}\dots x_{T}$. We first quantize the sequence into a sequence of indices $c_{1}\dots c_{T}$ with k-means clustering. In particular, $c_{t}=\operatorname*{arg\,min}_{i=1,\dots,k}\lVert x_{t}-m_{i}\rVert_{2}$, where $k$ is the number of clusters used in k-means clustering, and $m_{i}$ is $i$-th cluster centroid. The self-supervised loss function for training MelHuBERT is formulated as:
@@ II-A MelHuBERT
where
$p(c_{t}|x)$ $=\frac{\exp(W_{c_{t}}^{\top}o_{t})}{\sum_{i=1}^{K}\exp(W_{i}^{\top}o_{t})}$ (2)
is the distribution of the centroids predicted by the Transformer, $W$ is a learnable projection matrix, $W_{i}$ is the $i$-th row of $W$, and $o_{t}$ is the model’s output representation at time step $t$.
Following [28], we train two variants, MelHuBERT-10ms and MelHuBERT-20ms. MelHuBERT-10ms takes the typical Mel spectrograms at a 10 ms frame period, while MelHuBERT-20ms takes the concatenation of every two frames to achieve a 20 ms frame period. For MelHuBERT-20ms, we select the cluster labels of odd Mel frames $c_{1},c_{3},\dots,c_{2n+1}$, where $n=\lceil\frac{T}{2}\rceil-1$, as predicting targets. Similar to [3], we train MelHuBERT in two stages. The first stage uses quantized Mel spectrograms as targets, while the second stage uses quantized MelHuBERT hidden vectors as targets. Note that in the second stage, the frame rate of hidden vectors is the same as the targets, so we do not subsample the targets as in the first stage for MelHuBERT-20ms.
MelHuBERT not only is competitive against HuBERT but also trains at a lower computational cost, offering a strong footing for subsequent compression.
II-B Weight pruning
Weight pruning is one of the earliest pruning algorithms [29]. The idea is based on the finding that the increase in loss due to small changes to the weights of neural networks can be quickly recovered by a small amount of training. Weight pruning usually is an instance of iterative pruning, which converges faster and with lower training loss compared to one-shot pruning. In this work, we adopt one of the simplest criterion for pruning [30, 7], setting the weight to zero when the absolute value is below a threshold. We follow the usual implementation [7], using a binary mask to keep track of what weights are pruned. Weights that are pruned are never instantiated again.
To be more precise, suppose $f$ is the forward function, $x$ is the input, and $\theta$ is the model parameter. To compute the forward results of a pruned model, instead of computing $f(x;\theta)$, we compute $f(x;m\odot\theta)$, where $m\in\{0,1\}^{|\theta|}$ is a binary mask. The value of the mask $m$ is determined by:
@@ II-B Weight pruning
where $\delta$ is the threshold determined based on a target density. We will have a schedule for the amounts to be pruned in the experiments.
Since the mask does not exhibit any structure, speed-up of weight pruning can only come from sparse parallel computations or dedicated hardware [6]. In this work, we do not explore these options and only report theoretical speed-up.
II-C Head pruning
Head pruning is an iterative pruning technique specific to Transformers [8], where the unit of a block being pruned is a set of key, query, and value weights for computing self-attention. Previous studies have discovered that it is sufficient to perform well with only few attention heads [32, 33]. We are interested in to what extent this statement holds for self-supervised Transformers. If it holds, pruning heads can lead to significant reduction in memory and computation, considering the $O(T^{2})$ memory and computation for a sequence of length $T$ to compute the self-attention.
Similar to weight pruning, we compute a score for each head and prune heads of low scores to a target density. There are several options to compute the score of an attention head. The first is simply using the $\ell_{1}$ norm of the parameters. Formally, consider $N$ self-attention heads $H_{1},\dots,H_{N}$. For each head $H_{i}$, we compute a score
$s_{i}=\left\|\text{vec}(W_{i}^{Q})\right\|_{1}+\left\|\text{vec}(W_{i}^{K})\right\|_{1}+\left\|\text{vec}(W_{i}^{V})\right\|_{1},$ (4)
where $W_{i}^{Q}$, $W_{i}^{K}$, and $W_{i}^{V}$ are the weights of the query, key, and value matrices in $H_{i}$ respectively, and $\text{vec}(W)$ is a vector by flattening a matrix $W$. Another approach proposed in [33] is to use a small data set $\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ to compute the score
@@ II-C Head pruning
for each head $H_{i}$, where $J_{i}$ is the attention map multiplied by the forward result of the value matrix $V_{i}$, and $\frac{\partial\mathcal{L}}{\partial J_{i}}$ is the gradient to that particular computation node. This quantity considers the $\ell_{1}$ of both the attention map and its gradient. If the $\ell_{1}$ norm of the forward result is small, then the head is a good candidate to prune. If the $\ell_{1}$ norm of the gradient is small, changing the attention map has little effect on the loss, also meaning that the head is a good candidate to prune.
Contrary to weight pruning, the computation of self-attention can be avoided entirely when pruned, so the approach can lead to actual runtime speed-up without sparse operations or dedicated hardware.
II-D Low-rank approximation
Weight matrices in fully connected layers often turn out to be low-rank after training [9]. In this case, the low-rank matrices can be approximated by matrix factorization, for example, with singular value decomposition [34]. Transformers have two feed-forward layers after self-attentions, and the dimension after the first feed-forward layer is typically large (e.g., 3072). Most of the computation is due to the large dimension of this hidden layer. Given the high dimensionality, we suspect the weight matrices before and after the 3072-dimensional vector are low-rank. Instead of doing matrix factorization, we prune the 3072-dimensional hidden layers. This amounts to pruning the columns of the first feed-forward layer and the corresponding rows of the second feed-forward layer based on the sum of their weight magnitudes.
Formally, assume that $U\in\mathbb{R}^{d\times 3072}$ and $V\in\mathbb{R}^{3072\times d}$ to be the projection matrix of the first feed-forward layer and the second feed-forward layer of a Transformers layer respectively, where $d$ is the hidden dimension of the model. For each dimension $i$, we compute a score
$s_{i}=\sum_{k=1}^{d}\left(|U_{ki}|+|V_{ik}|\right).$ (6)
We then prune the rows and columns of low scores to a target density. If $U$ and $V$ are low-rank, pruning the columns of $U$ and the rows of $V$ would likely not reduce their rank.
Note that our approach differs from the usual low-rank approximation. Low-rank approximation as used in [9] maintains the input and output dimensions, and offers speed-up by factorizing one large matrix into two matrices with a low-dimension bottleneck. In contrast, our approach involves two matrices, and offers speed-up by pruning rows and columns.
II-E Knowledge distillation
Knowledge distillation is often used to compress a model, training a smaller neural network (a student) to match parts of the other, larger neural network (a teacher) [35]. Matching the output of a teacher is the most common, while matching the hidden layers are also a viable option [36, 37, 38]. In terms of the losses, KL divergence is often used to match probability outputs, while $\ell_{2}$ norm is often used to match the hidden vectors [38]. In this work, since MelHuBERT uses cross entropy for training, we simply use KL divergence as the distillation loss. Specifically, suppose the output representation of the student at the timestamp $t$ is $o_{t}$. We minimize the loss
@@ II-E Knowledge distillation
Knowledge distillation is fundamentally different from other pruning approaches, in that a small, student model is trained from random initialization. This approach does not attempt to preserve the structure of the teacher model, nor does it preserve the weights of the teacher model. This approach does not even rely on the training loss that the teacher model is trained on; hence the popularity among recent compression approaches, such as DistilHuBERT [19], FitHuBERT [21], and LightHuBERT [20].
III Related work
There are many attempts to compress self-supervised models in text and vision applications, and we broadly mention a few that are based on Transformers. Some of them [39, 40, 41, 42] apply knowledge distillation to transfer knowledge from the teacher model. Others [33, 43, 44] decide to prune parts of the network by either weight pruning or head pruning. Additionally, some [45] use the concept of low-rank approximation to factorize weight matrices. Some [46, 47, 48] choose to quantize the network to a low-precision version. Another way to speed up a model is to stop the forward process midway [49, 50]. Other than the low-precision approaches, these studies inspire our choice of approaches and design of experiments to be detailed in later sections.
In speech processing, many studies focus on producing a compressed model from a pre-trained self-supervised model for a particular downstream task. This approach differs from regular compression in that models are simultaneously fine-tuned for a particular task and compressed. Examples along this line include [14, 15, 16, 17].
Other studies that do not focus on a specific downstream task, instead, compress models at the pre-training stage. These approaches either require access to the pre-training loss (like what is done in this paper) or have to resort to knowledge distillation [19, 20, 21, 22, 23, 24, 25, 26, 27]. These types of compressed models are capable of extracting representations that can generalize to multiple downstream tasks.
DistilHuBERT [19] is one of the first attempts to study distillation on speech self-supervised model. They introduce three prediction heads after the final layer of the student model to match the hidden representations of the fourth, eighth, and twelfth layers of the teacher model. Cosine similarity and $\ell_{1}$ norm are combined as their loss function. They use the weight of the first two layers of the teacher model to initialize the student model. Their result, DistilHuBERT, is a 2-layer Transformer, demonstrating the viability of distilling a self-supervised speech model.
FitHuBERT [21] is also a model based on distillation. Compared to DistilHuBERT, they choose to keep the model depth at 12 layers while reducing the dimensions of self-attention and feedforward layers. Additionally, they decrease the number of channels in the lower layers of the convolutional feature extractor, significantly reducing the computational cost.
LightHuBERT [20] involves a two-stage training, where a HuBERT-sized supernet is trained by distillation in the first stage. Then, they apply once-for-all training in the second stage. In particular, they divide the Transformer into several parts, such as the embedding dimension, the number of attention heads, the width of the feedforward network, and the depth of the Transformer. They then randomly sample a subnet from the student model according to the predefined structure during each forwarding pass. Their smallest model, LightHuBERT small, is a 12-layer Transformers with reduced embedding dimension, reduced feedforward layer hidden dimension, and a reduced number of self-attention heads. LightHuBERT small achieves strong performance, but requires an additional distillation to a HuBERT-sized network, a stringent requirement without many industrial-scale GPUs.
Peng et al. [17] propose an approach to simultaneously fine-tuned a self-supervised model particularly for ASR while compressing heads and feedforward layers in Transformers. Wang et al. [25] extend their methods to a task-agnostic version by replacing the objective for ASR with knowledge distillation. Their compressed model is able to extract general representation across different downstream tasks, and can be seen as a combination of knowledge distillation and pruning.
In this work, we will compare our approach to DistilHuBERT, FitHuBERT, and LightHuBERT. We do not compare to Wang et al. [25] because of two reasons. They use knowledge distillation, while we use the pre-training loss for pruning, more faithful to the pruning objectives in the supervised setting. They prune both feedforward layers and attention heads together, confounding the sources of improvements. We, instead, offer a clean comparison among approaches and identify the their strengths and weaknesses. We also offer, as an application of our results, a simple approach to combining compression techniques.

### v3

# Is Smaller Always Faster? Tradeoffs in Compressing Self-Supervised Speech Transformers

@@ Abstract
Transformer-based self-supervised models have achieved remarkable success in speech processing, but their large size and high inference cost present significant challenges for real-world deployment. While numerous compression techniques have been proposed, inconsistent evaluation metrics make it difficult to compare their practical effectiveness. In this work, we conduct a comprehensive study of four common compression methods, including weight pruning, head pruning, low-rank approximation, and knowledge distillation on self-supervised speech Transformers. We evaluate each method under three key metrics: parameter count, multiply-accumulate operations, and real-time factor. Results show that each method offers distinct advantages. In addition, we contextualize recent compression techniques, comparing DistilHuBERT, FitHuBERT, LightHuBERT, ARMHuBERT, and STaRHuBERT under the same framework, offering practical guidance on compression for deployment.
@@
Speech self-supervised model, model compression
@@ I Introduction
Self-supervised learning (SSL) has driven significant progress in speech processing, enabling models to learn powerful representations from large unlabeled datasets [1, 2, 3, 4]. However, these models are typically large and compute-intensive, making them difficult to deploy in real-time or on resource-constrained devices. Model compression is therefore essential to bridge the gap between research prototypes and practical deployment.
Numerous techniques have been proposed for compressing SSL models, including weight pruning [5], low-rank approximation [6], and knowledge distillation [7]. Yet, these methods are often evaluated in isolation and compared under different metrics, such as parameter count, multiply-accumulate operations (MACs), or inference time. This inconsistency makes it difficult to assess their practical strengths and tradeoffs. A method that reduces parameter count might not improve actual inference speed, and one that accelerates GPU runtime might have little impact on theoretical complexity. Without a consistent evaluation framework, it remains unclear which method is best suited for which use case.
In this paper, we conduct a comprehensive study on compressing Transformer-based speech SSL models. We systematically compare four classic compression methods—weight pruning, head pruning, low-rank approximation, and knowledge distillation—under three critical metrics: number of parameters, MACs, and real-time factor.
Our results show that each method exhibits distinct strengths under different metrics: weight pruning significantly reduces parameter count and MACs; head pruning improves inference speed by reducing attention computations; low-rank approximation effectively reduces parameter counts; and distillation provides the most balanced tradeoff between efficiency and performance.
To further clarify the efficiency-performance tradeoffs in recent research, we also contextualize several representative compression methods, including DistilHuBERT [8], FitHuBERT [9], LightHuBERT [10], ARMHuBERT [11], and STaRHuBERT [12] within the same framework. This unified comparison not only highlights how design decisions impact compression outcomes, but also offers practical guidance for selecting methods based on deployment constraints.
II Related Work
Model compression has been widely studied in the context of NLP and computer vision, particularly for Transformer-based architectures. Common strategies include knowledge distillation [7, 13], weight pruning [14], head pruning [15], low-rank approximation [6], and early exiting [16, 17]. In the speech domain, early efforts on compressing SSL models have largely focused on task-specific settings, where compression is jointly optimized with downstream objectives [18, 19, 20, 21].
Recent studies have proposed task-agnostic compression methods that aim to preserve the general usability of speech representations. DistilHuBERT [8] distills a 12-layer HuBERT into a shallow 2-layer student by aligning hidden representations using cosine and $\ell_{1}$ losses. FitHuBERT [9] retains the model depth but reduces the width of attention heads, feedforward layers, and convolutional channels to achieve a thinner architecture. LightHuBERT [10] introduces a two-stage training process: first distilling a large model into a supernet, then applying once-for-all training to support a range of subnetwork configurations. ARMHuBERT [11] and STaRHuBERT [12] share a similar architecture with FitHuBERT but achieve superior performance through masking distillation and temporal relation distillation, respectively. Beyond distillation-based approaches, several task-agnostic, pruning-based methods have been proposed for speech self-supervised models [22, 23]. Additionally, Lin et al. [24] propose a method that dynamically determines whether to perform early exiting during feature extraction.
While these prior works report promising results, they each adopt different evaluation metrics—DistilHuBERT, LightHuBERT, and ARMHuBERT report only parameter count; FitHuBERT reports parameter count and real-time factor; and STaRHuBERT reports parameter count and MACs—making it difficult to directly compare their efficiency-performance tradeoffs. Our unified framework addresses this gap by providing consistent and comprehensive evaluations.
III Model Compression
We plan to study four classic compression methods, weight pruning, head pruning, low-rank approximation, and knowledge distillation. All except distillation are instances of iterative pruning, where pruning and retraining alternate until convergence. We intentionally exclude more involved compression strategies that integrate task objectives or joint pruning components in order to isolate and clarify the effect of each technique.
After pruning, model performance typically degrades due to increased training loss. To recover performance, additional training is applied. To track recovery, we monitor the self-supervised loss curve before and after compression. We adopt MelHuBERT [25] as our base model for its simple training recipe, which allows efficient pre-training from scratch and loss monitoring.
III-A MelHuBERT
MelHuBERT simplifies HuBERT by using Mel spectrograms as input and cross-entropy as the training loss. It removes waveform-level convolutions and accelerates training. MelHuBERT saves about 33.3% of MACs and about 31.2% of pre-training time on a single 24GB RTX 3090 GPU. Given a Mel-spectrogram sequence $x_{1}\dots x_{T}$, we quantize each frame into cluster indices $c_{t}$ using k-means. The training objective is:
@@ III-A MelHuBERT
$\quad p(c_{t}|x)=\frac{\exp(W_{c_{t}}^{\top}o_{t})}{\sum_{i=1}^{K}\exp(W_{i}^{\top}o_{t})},$ (2)
where $W$ is the classifier matrix and $o_{t}$ the model output. This loss is used throughout our compression analysis.
III-B Weight Pruning
Weight pruning is one of the earliest model compression techniques [26], based on the observation that small weights can be removed with minimal impact on performance, and retraining can recover any loss. In iterative pruning, we progressively remove low-magnitude weights and fine-tune the model. We adopt a simple magnitude-based criterion [27, 28], where weights are masked as follows:
@@ III-B Weight Pruning
where $\delta$ is a threshold set by the target sparsity. The forward pass becomes $f(x;m\odot\theta)$, where $m$ is the binary mask. While weight pruning significantly reduces parameters and MACs, the unstructured sparsity limits real-time gains on standard hardware.
III-C Head Pruning
Head pruning targets Transformer-specific structure by removing full attention heads, each comprising its own query, key, and value matrices. Prior work [29, 15] shows that many heads are redundant. We evaluate the impact of pruning heads in speech SSL models. While several methods have been proposed, we adopt a gradient-based importance score from [15], since simpler $\ell_{1}$-norm-based scores perform consistently worse in our experiments. Specifically, we use a small data set $\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ to compute the score
@@ III-C Head Pruning
for each head $H_{i}$, where $J_{i}$ is the attention map multiplied by the forward result of the value matrix $V_{i}$, and $\frac{\partial\mathcal{L}}{\partial J_{i}}$ is the gradient to that particular computation node.
III-D Low-Rank Approximation
Low-rank approximation leverages the empirical low-rank nature of learned weight matrices [30]. Rather than decomposing projection matrices, we prune intermediate dimensions in feed-forward layers, where the hidden size is typically large (e.g., 3072). Let $U\in\mathbb{R}^{d\times 3072}$ and $V\in\mathbb{R}^{3072\times d}$ be the input/output projections. We compute a score for each hidden unit $i$ as:
$s_{i}=\sum_{k=1}^{d}(|U_{ki}|+|V_{ik}|).$ (5)
We iteratively remove units with the lowest scores. Unlike conventional low-rank factorization, this direct pruning approach simplifies implementation while achieving comparable efficiency.
III-E Knowledge Distillation
Knowledge distillation compresses models by training a small student to imitate a large teacher [31]. It allows full architectural flexibility. We distill from a 12-layer MelHuBERT into 2- and 6-layer students using KL divergence on output distributions:
@@ III-E Knowledge Distillation
Unlike pruning-based methods, distillation trains a new model from scratch and does not preserve teacher weights or structure. It is widely used in prior SSL compression works [8, 9, 10].